OpenStudy (annipuppi):

1. When solving a system of equations how do you know which method to use? 2. Solve the system of equations below and explain the method you used and show the steps you used. –y + 3x = 6 y = –6x + 12 Please explain it because I'm really confused on this.

OpenStudy (amistre64):

the method you use is the one that you are most comfortable with .... method is only for user functionality

OpenStudy (amistre64):

spose we use elimination –y + 3x = 6 y = 12 - 6x ----------------- 3x = 18 - 6x ; solve for x, to determine y if we use substitution –y + 3x = 6 y = 12 - 6x <-- it is already known that y = some function of x so sub it into the other one –(12-6x) + 3x = 6 ; and soplve for x to determine the value of y

OpenStudy (annipuppi):

I don't really understand it still. can you try and explain how to do solve the equation better?

OpenStudy (anonymous):

R1: 3x-y=6 R2: 6x+y=12 solve this by using elimination method R2+R1 => 9x = 18 => x=2 and then replace x=2 into R1 or R2 and solve for y

OpenStudy (annipuppi):

Oh my gosh guys I'm so confused! I'm having a lot of trouble in math right now. Can you like label each step and expalin it? Please. I know I probably sound stupid since i cant fiqure this out when you gave me the answer.

OpenStudy (amistre64):

are you comfortable with being able to solve for a single variable?

OpenStudy (annipuppi):

Yeah liek yo mean in a normal equation?

OpenStudy (amistre64):

yep

OpenStudy (amistre64):

the idea is to reduce this setup to a single variable that can be solved; and once you know the value of one of the variables, that can be used to determine the value of the other one

OpenStudy (amistre64):

the method of substitution is simply defining one variable in terms of the other; notice that one equation they give you already defines y in terms of x y = –6x + 12 ; so no matter what y equals at the moment, it is completely dependant on the value of x; whereever we see a y, we can replace it by the equation that defines it; does that make sense?

OpenStudy (annipuppi):

Yeah

OpenStudy (amistre64):

so lets take a look at the other equation they provide; and see if there is a y variable in it that we can redefine by its equation in x –y + 3x = 6 ^ theres a y, lets replace it –(-6x+12) + 3x = 6 ; now we have a setup that we can solve for a single variable right?

OpenStudy (annipuppi):

Yeah it makes a lot more sense now.

OpenStudy (amistre64):

once we have a solid value for x; we can use it to define y :) since: y = -6x+12 right?

OpenStudy (annipuppi):

Yeah so you just solve the one equation to find x then you solve y=-6+13 to find y and then you right the answer like (x,y) right?

OpenStudy (amistre64):

exactly ;)

OpenStudy (annipuppi):

Okay thanks. I get it now.

OpenStudy (amistre64):

youre welcome, and good luck ;)